On discovering that her family had a 70% risk of heart attack, Erin took a treadmill test to check her own potential of having a heart attack. The doctors told her that the reliability of the stress test is 67%. The test predicted that Erin will not have a heart attack. What is the probability after the test was taken that she will have a heart attack?


0.4051

0.5010

0.4653

0.6632

Verified Answer :

Answer:

C.  0.4653

Step-by-step explanation:

This a conditional probability problem where Bayes theorem should be applied.

Applying Bayes theorem,

\(P(\text{No heart attack}\ |\ \text{Correctly tested})=\)

\(\dfrac{P(\text{Correctly tested}\ |\ \text{No heart attack})\cdot P(\text{No heart attack})}{P(\text{Correctly tested})}\)

\(P(\text{Correctly\ tested}\ |\ \text{No\ heart\ attack})=67\%=0.67\)

\(P(\text{No\ heart\ attack})=1-P(\text{heart\ attack})=1-0.7=0.3\)

\(P(\text{Correctly\ tested})=[P(\text{No\ heart\ attack})\times P(\text{Correctly\ tested})]+[P(\text{Heart\ attack})\times (\text{Incorrectly\ tested})]\)

\(=[0.3\times 0.67]+[0.7\times 0.33]=0.432\)

Putting the values,

\(P(\text{No\ heart\ attack}\ |\ \text{Correctly\ tested})=\dfrac{0.67\times 0.3}{0.432} =0.4653\)

There is a probability of 0.4653 or 46.53% chance that she will not have a heart attack even though the test predicts that she will.

Answer:

C= .4653

Step-by-step explanation:

i just took the test and got it right